OpenStudy (ashleyy_smith):

Use substitution to solve the system. −2x+3y=14 x-2y=−8

rebeccaxhawaii (rebeccaxhawaii):

okay so first find out what y is from one equation

rishavraj (rishavraj):

@Ashleyy_Smith really ???

OpenStudy (razork):

Are these 2 seperate problems?

OpenStudy (ashleyy_smith):

I was trying to do it on paper but it was hard for me I got stuck

rishavraj (rishavraj):

look at the 2nd equation u got the value of x in terms of y .... so u can substitute is easily @Ashleyy_Smith

rishavraj (rishavraj):

*substitute it easily in the 1st equation

rishavraj (rishavraj):

@Ashleyy_Smith ????

OpenStudy (ashleyy_smith):

Is it 4 for the first equation?

rishavraj (rishavraj):

see re-arrange the 2nd equation so it would be x = 2y - 8 u get that

OpenStudy (ashleyy_smith):

no that's the thing I don't get

rishavraj (rishavraj):

x - 2y = -8 x -2y + 2y = -8 + 2y

OpenStudy (ashleyy_smith):

ohhhhh

rishavraj (rishavraj):

now u know wht to do/?

rishavraj (rishavraj):

see u can even use equation 1 to get value of any one variable in terms of the other....bt tht would be really hectic :))

OpenStudy (ashleyy_smith):

So I just do what I did with the first equation right?

rishavraj (rishavraj):

hmmm wht do u mean???

rishavraj (rishavraj):

see u got x = 2y -8 now plug the value of x in 1st equation...as did the previos time

OpenStudy (ashleyy_smith):

-2x+3y=14 -2(2y-8)+3 = 14

rishavraj (rishavraj):

y missing >_<

rishavraj (rishavraj):

hmmm ??? :/

OpenStudy (ashleyy_smith):

oh oops after the 3 :p

OpenStudy (ashleyy_smith):

-4y + 16 + 3y = 14

rishavraj (rishavraj):

go on :))

OpenStudy (ashleyy_smith):

-7y +16 = 14?

rishavraj (rishavraj):

hold on aint u kiddinng me -4y + 3y = -7y ???? -_-

OpenStudy (ashleyy_smith):

oh oops y

OpenStudy (ashleyy_smith):

forgot there was a negative

rishavraj (rishavraj):

really y?????

rishavraj (rishavraj):

whts 4 - 5??

OpenStudy (ashleyy_smith):

-1....... lol

rishavraj (rishavraj):

so -4y + 3y = ??

OpenStudy (ashleyy_smith):

-1 :p I thought it would be y cause I thought y is considered like 1

rishavraj (rishavraj):

-y -_-

OpenStudy (ashleyy_smith):

yes

rishavraj (rishavraj):

-4 y + 16 + 3y = 14 -y + 16 = 14 now wht

OpenStudy (ashleyy_smith):

hmm i dont know

rishavraj (rishavraj):

isolate the variable on one side...... like wht would be x , if x + 2 = 5??

OpenStudy (ashleyy_smith):

3

rishavraj (rishavraj):

so whts y , if -y + 16 = 14

OpenStudy (ashleyy_smith):

0

OpenStudy (ashleyy_smith):

lol

rishavraj (rishavraj):

how?? :/

OpenStudy (ashleyy_smith):

-2 jk lol. wanted to make u sad

rishavraj (rishavraj):

so its -y = -2 so y = ??? btw instead of sad I get pissed off -_-

OpenStudy (ashleyy_smith):

thanks for letting me know ;) bahahah

OpenStudy (ashleyy_smith):

won't do it anymore. maybe.

rishavraj (rishavraj):

ok .... :)) y = ???

OpenStudy (ashleyy_smith):

-2

rishavraj (rishavraj):

why so ??? see -y = -2 y = ???

OpenStudy (ashleyy_smith):

2

rishavraj (rishavraj):

did u get tht ??? why y = 2 ?

OpenStudy (ashleyy_smith):

negative and negative = positive

rishavraj (rishavraj):

makes a bit sense.... :))) so as y = 2 wht would be x ??

OpenStudy (ashleyy_smith):

-3

rishavraj (rishavraj):

dazzle me -_- how ???

OpenStudy (ashleyy_smith):

3 lol

rishavraj (rishavraj):

and howz tht ???

rishavraj (rishavraj):

really I am jst asking to plug the value of 'y' in any of the 2 equations and get x.... so get me 'x'

OpenStudy (ashleyy_smith):

ohhhhh

OpenStudy (ashleyy_smith):

i think 2nd equation would be simple right

rishavraj (rishavraj):

yup

OpenStudy (ashleyy_smith):

so x-2(2) = -8

rishavraj (rishavraj):

go on.....

OpenStudy (ashleyy_smith):

x-4 = -8

rishavraj (rishavraj):

go on....

OpenStudy (ashleyy_smith):

x - 4 = -8 - 4 -4 x = -12

rishavraj (rishavraj):

How could u do this..... >_> see u need to isolate x -4 -4 = -8 -_-

OpenStudy (ashleyy_smith):

oh oops -4

rishavraj (rishavraj):

x = -4 ???

OpenStudy (ashleyy_smith):

no 2

rishavraj (rishavraj):

why ??? rhough x = -4 was correct -_-

rishavraj (rishavraj):

*though

OpenStudy (ashleyy_smith):

oh thought we had to divide the -4 and -8

OpenStudy (ashleyy_smith):

so x = -4 and y = 2

rishavraj (rishavraj):

why divide ??// see when x - 2 = 5 u add 2 on both sides

rishavraj (rishavraj):

and if 3x = 2 in this case we divide by 3 /.... :/

rishavraj (rishavraj):

and yeah x = -4 and y = 2

OpenStudy (ashleyy_smith):

yay finally

rishavraj (rishavraj):

listen it took another 1hour today.... Do me a favor and dont let it go worthless ....

rishavraj (rishavraj):

ok ??? @Ashleyy_Smith

OpenStudy (ashleyy_smith):

yes

OpenStudy (ashleyy_smith):

@rishavraj next topic will be elimination

rishavraj (rishavraj):

lol k...bt dont forget to practice this.....

OpenStudy (ashleyy_smith):

I won't thank u (:

OpenStudy (ashleyy_smith):

goodluck on your test

rishavraj (rishavraj):

u r most welcome :)))

rishavraj (rishavraj):

ooo thanks ...I need tht ...I dont even know the syllabus.... or which test I gotta take in the first place @Ashleyy_Smith

OpenStudy (ashleyy_smith):

:o! that's always the worst. Well, I'm going to eat dinner. I'm very hungry. I'll be on tomorrow (: thanks again!

rishavraj (rishavraj):

Bye :))